# Empowering Agentic Video Analytics Systems with Video Language Models

### v3

@@ Abstract.
AI-driven video analytics has become increasingly important across diverse domains. However, existing systems are often constrained to specific, predefined tasks, limiting their adaptability in open-ended analytical scenarios. The recent emergence of Video-Language Models (VLMs) as transformative technologies offers significant potential for enabling open-ended video understanding, reasoning, and analytics. Nevertheless, their limited context windows present challenges when processing ultra-long video content, which is prevalent in real-world applications. To address this, we introduce Avas, a VLM-powered system designed for open-ended, advanced video analytics. Avas incorporates two key innovations: (1) the near real-time construction of Event Knowledge Graphs (EKGs) for efficient indexing of long or continuous video streams, and (2) an agentic retrieval-generation mechanism that leverages EKGs to handle complex and diverse queries. Comprehensive evaluations on public benchmarks, LVBench and VideoMME-Long, demonstrate that Avas achieves state-of-the-art performance, attaining 62.3% and 64.1% accuracy, respectively, significantly surpassing existing VLM and video Retrieval-Augmented Generation (RAG) systems. Furthermore, to evaluate video analytics in ultra-long and open-world video scenarios, we introduce a new benchmark, Avas-100. This benchmark comprises 8 videos, each exceeding 10 hours in duration, along with 120 manually annotated, diverse, and complex question-answer pairs. On Avas-100, Avas achieves top-tier performance with an accuracy of 75.8%.
The Work is done during internship at Microsoft Research
copyright: acmcopyrightcopyright: rightsretainedconference: Preprint; price: ;
1. Introduction
@@
Video analytics (Jiang et al., 2021a; Bhardwaj et al., 2022; Khani et al., 2023; Nan et al., 2023) has emerged as a transformative technology across a wide array of domains, such as security surveillance, intelligent transportation systems, industrial automation, and retail monitoring. By harnessing the power of deep learning models, video analytics systems extract meaningful insights, identify patterns, and generate actionable information from video data, enabling organizations to monitor, detect, and respond to events with greater efficiency and precision. The desired features of video analytics systems necessitate a variety of capabilities. We classify the intelligence levels of existing and future video analytics systems into five levels, labeled as L1 through L5:
@@
L1 systems for specific classification, segmentation, and detection using models e.g., ResNet (He et al., 2016) and EfficientDet (Tan et al., 2020) to extract spatial information from video data, including object classes and bounding boxes, etc (Jiang et al., 2021a; Bhardwaj et al., 2022; Khani et al., 2023).
@@
L2 systems go beyond spatial information extraction by enabling causal event detection and analytics, i.e., identifying short-term events. They use models like C3D (Tran et al., 2015) and ActionFormer (Zhang et al., 2022) to detect and localize events (e.g., actions, activities, anomalies) through spatiotemporal modeling.
L3 systems extend L2’s spatiotemporal detection by incorporating neural language processing (NLP). Using models like CLIPBERT (Lei et al., 2021), they interpret and respond to natural language queries (e.g., ”What animals appeared in the videos?”) instead of merely outputting fixed labels. Though queries are still limited to specific domains, L3 systems greatly improve user interaction for video analytics.
@@
Despite significant advancements, current video analytics systems (Khani et al., 2023; Bhardwaj et al., 2022; Zhang et al., 2024b; Padmanabhan et al., 2023; Lu et al., 2022; Sivaraman et al., 2024; Zhang et al., 2021; Jiang et al., 2021a) primarily focus on L1 to L3 levels. These systems are designed for closed-end analytics, relying on specialized models for specific tasks, which limits their flexibility and adaptability. Consequently, we envision L4 and L5 systems to enable open-end analytics.
@@
L4 systems enable open-ended video comprehension, reasoning, and analytics, marking a major advancement. They support general-purpose analytics on video sources, handling natural language queries and responses while enabling complex, long-term spatiotemporal reasoning. For example, they can answer questions like: ”What abnormal events occurred in the past ten hours?”, ”What caused the person to fall?”, or ”How did the animals behave after appearing on camera?”
@@
L5 systems advance L4 by autonomously integrating external public or domain-specific knowledge to uncover explicit and implicit connections between videos and world knowledge, fulfilling the ultimate goal of video analytics: deriving meaningful insights and offering actionable automations.
@@
In this paper, we delve into the development of L4 video analytics systems, leveraging the transformative potential of video-language models (VLMs). By combining vision and language understanding, VLMs enable generalized visual detection and advanced video comprehension, including causal reasoning, key information retrieval, and human-interpretable explanations. The integration of perception, reasoning, and interaction makes VLMs highly adaptable, positioning them as a key foundation for L4 systems.
@@
However, integrating VLMs into video analytics poses significant challenges, primarily due to the limited context window of current VLMs compared to the extensive duration of video sources in typical video analytics scenarios. While L1 to L3 systems handle spatial and short-term causal event detection by processing frames independently or within small sliding windows (e.g., a few seconds), L4 systems require collective analysis of related frames for long-term causal detection, summarization, and reasoning. Current VLMs, like QwenVL (Wang et al., 2024a), can processing up to 768 frames, covering minutes or hours of video. However, video analytics often involves much longer sources, spanning hundreds of hours or continuous streams, far beyond the capabilities of existing VLMs.
@@
Recent studies (Shen et al., 2024, 2025; Wang et al., 2025) that attempt to extend the context window of VLMs remain inadequate for processing video sources spanning hundreds of hours. Retrieval-augmented generation (RAG) frameworks (Edge et al., 2024; Guo et al., 2024; Fan et al., 2025; Research, 2024) aim to address similar limitations by first retrieving relevant frames from massive contents and then generating final answers. However, these approaches still face significant challenges in handling the video modality (as detailed in §7), leading to both reduced analytics accuracy and substantial computational overhead.
@@
In this paper, we present Avas, a system that integrates VLMs into video analytics to enable L4 capabilities. The core innovation of Avas is its efficient indexing mechanism, designed to handle extremely long video sources or unlimited video streams, and by leveraging the index, Avas effectively retrieves relevant information for a given query and generates accurate, robust responses. Specifically, Avas introduces the following key features that distinguish it from existing systems: 1) analyzing extremely long videos, spanning hundreds of hours or even unlimited video streams; 2) supporting for near-real-time (e.g., at more than 1 FPS) index construction and analytics; and 3) handling diverse and complex queries including temporal grounding and reasoning, summarization, event and entity understanding, and key information retrieval.
@@
Particularly, Avas introduces two key components: near-real-time index construction and agentic retrieval and generation. During the index construction phase (§4), we propose event knowledge graphs (EKGs) as an indexing mechanism for video analytics. Unlike traditional knowledge graphs (KGs) used in text-based RAG systems (Edge et al., 2024; Guo et al., 2024; Research, 2024), EKGs represent a flow of insightful events, effectively capturing video dynamics and temporal consistency. Using a small VLM, such as QwenVL-7B, Avas extracts information to construct the EKG. To optimize this process, we introduce techniques that enable near-real-time index construction, achieving more than 5 FPS on typical edge servers equipped with 2 $\times$ 4090 GPUs.
@@
In the retrieval and generation phase (§5), Avas employs an agentic search mechanism instead of directly retrieving information from the constructed index. This approach allows Avas to proactively retrieve more relevant information by utilizing contextual hints captured within the EKG, enables Avas to handle complex queries, including summarization and multi-hop reasoning etc. Furthermore, we introduce techniques to enhance Avas to robustly generate answers based on the retrieved information.
@@
We evaluate Avas on multiple public video understanding benchmarks, including LVBench (Wang et al., 2024b) and VideoMME-Long (Fu et al., 2024). These benchmarks collectively comprise approximately 400 videos and 2,500 questions. We compare Avas against a diverse range of baselines, including state-of-the-art (SOTA) VLMs such as GPT-4o (Achiam et al., 2023), Gemini (Team et al., 2023), Phi-4-Multimodal (Abdin et al., 2024a), Qwen2.5-VL-7B (Wang et al., 2024a), InternVL2.5-8B (Chen et al., 2024), and LLaVA-Video-7B (Zhang et al., 2024a), as well as typical video-RAG systems such as VideoTree (Wang et al., 2024c), VideoAgent (Wang et al., 2024d), DrVideo (Ma et al., 2024), and VCA (Yang et al., 2024). On both benchmarks, Avas establishes new SOTA performance, achieving 62.3% on LVBench and 64.1% on VideoMME-Long, respectively, significantly surpassing baselines by up to 16.9% and 5.2%.
@@
In addition to the public video understanding benchmarks, we introduce a new benchmark, Avas-100, specifically designed to evaluate L4 video analytics tasks. Avas-100 comprises 8 ultra-long videos, each exceeding 10 hours in duration, and includes a total of 120 manually annotated questions and answers. The videos are carefully curated from typical video analytics scenarios, and the questions cover multiple key categories relevant to video analytics systems. Evaluation results show that Avas achieves significantly better performance on Avas-100 compared to various baselines, with improvements of approximately 20.8%. In summary, we make the following contributions in this paper:
@@
  • We propose Avas, the first L4 video analytics system powered by VLMs, to the best of our knowledge.
@@
  • Avas introduces near-real-time index construction and agentic retrieval and generation, along with innovative techniques that enable key features for L4 video analytics, including open-ended analytics on extremely long video sources in near-real-time.
@@
  • We evaluate Avas on two widely-used video understanding benchmarks, achieving SOTA performance with 62.3% on LVBench and 64.1% on VideoMME-Long.
@@
  • Furthermore, we present Avas-100, a benchmark specifically designed for L4 video analytics systems, where Avas demonstrates significant improvements, outperforming baselines by approximately 20.8%.
2. Related Work and Motivation
2.1. Video Analytics System and VLMs
@@
The field of video analytics (VA) has seen significant advancements in recent years (Khani et al., 2023; Bhardwaj et al., 2022; Zhang et al., 2024b; Padmanabhan et al., 2023; Lu et al., 2022; Sivaraman et al., 2024; Zhang et al., 2021). Leveraging emerging deep learning models, modern VA systems can extract insightful information, such as object locations or counts, from video streams processed on edge or cloud servers.
@@
Existing VA systems predominantly support closed-end analytics (L1 to L3 systems as mentioed in §1), often relying on shallow neural networks to extract predefined, task-specific, and constrained information. For instance, Remix (Jiang et al., 2021a) leverages fine-tuned EfficientDet (Jiang et al., 2021b) to generate bounding boxes for pedestrians. Consequently, the flexibility and adaptability of VA systems are fundamentally limited by the inherent constraints of the specific models they employ.
@@
Recently, video language models (VLMs), e.g., GPT-4o (Achiam et al., 2023), Gemini (Team et al., 2023, 2024), QwenVL (Wang et al., 2024a; Bai et al., 2025) and Phi (Abdin et al., 2024a, b; Abouelenin et al., 2025), have demonstrated their transformative potential in video analytics tasks. By leveraging the extensive world knowledge embedded in large language models (LLMs), VLMs not only achieve generalized visual grounding but also exhibit advanced video comprehension capabilities, such as zero-shot temporal and spatial reasoning, contextual retrieval, and semantic understanding. More importantly, VLMs enable natural language interaction, allowing users to dynamically query video content. This makes them particularly well-suited for addressing diverse and unstructured open-ended analytics in real-world scenarios, towards L4 VA systems.
@@
However, adopting VLMs in VA systems is far from straightforward. Existing L1 to L3 VA systems, which rely on traditional DNNs, typically process each video frame independently. In contrast, VLMs require related frames to be processed collectively to infer causal relationships and temporal dependencies across frames. This shift introduces significant complexity, as existing VLMs are generally capable of handling only minute-level or sub-hour-level videos due to the limited context window inherent in language models.
@@
In real-world video analytics scenarios, the scale of videos to be analyzed is often vastly larger—spanning hundreds of hours or more (e.g., monitoring wildlife behavior over an entire month, as illustrated in Fig. 2). This creates a fundamental gap between the capabilities of VLMs and the demands of VA systems, as the limited context window length of VLMs directly restricts their ability to process videos of such extensive durations effectively.
2.2. Long Video Understandings
Recent research also increasingly focuses on enabling long video understandings (Shen et al., 2024; Wang et al., 2025). Given the inherent limitations of autoregressive language models, such as the constrained context window length, efforts have been directed toward reducing the redundancy in video inputs to facilitate the processing of extended video durations. For instance, LongVU (Shen et al., 2024) and AdaRETAKE (Wang et al., 2025) introduce dynamic compression mechanisms that prioritize video content based on its relevance, selectively retaining frames or regions most pertinent to downstream language tasks. Similarly, NVILA (Liu et al., 2024) addresses the efficiency-accuracy trade-off by optimizing sampling strategies and resolution to fit within limited token budgets.
While these approaches have succeeded in increasing the number of frames that models can process and mitigating the constraints of context windows to some extent, they fall short of achieving a fundamental breakthrough. Most existing methods remain limited, with the supported video length typically capped at approximately one hour (Qian et al., 2024), which is insufficient to meet the demands of video analytics. Moreover, as video length extends, the inference cost rises correspondingly, further exacerbating the challenges of scaling these systems.
2.3. Retrieval Augmented Generation
Figure 1. The system overview of Avas.
Short (1.4 minutes) Medium (9.7 minutes) Long (39.7 mintues)
@@ 2.3. Retrieval Augmented Generation
Table 1. A small portion of frames is required to answer each specific question across the short, medium, and long video subsets of the VideoMME (Fu et al., 2024) benchmark using Qwen2-VL.

### v5

# Ava: Towards Agentic Video Analytics with Vision Language Models

@@ Abstract
AI-driven video analytics has become increasingly important across diverse domains. However, existing systems are often constrained to specific, predefined tasks, limiting their adaptability in open-ended analytical scenarios. The recent emergence of Vision Language Models (VLMs) as transformative technologies offers significant potential for enabling open-ended video understanding, reasoning, and analytics. Nevertheless, their limited context windows present challenges when processing ultra-long video content, which is prevalent in real-world applications. To address this, we introduce Ava, a VLM-powered system designed for open-ended, advanced video analytics. Ava incorporates two key innovations: (1) the near real-time construction of Event Knowledge Graphs (EKGs) for efficient indexing of long or continuous video streams, and (2) an agentic retrieval-generation mechanism that leverages EKGs to handle complex and diverse queries. Comprehensive evaluations on public benchmarks, LVBench and VideoMME-Long, demonstrate that Ava achieves state-of-the-art performance, attaining 62.3% and 64.1% accuracy, respectively, significantly surpassing existing VLM and video Retrieval-Augmented Generation (RAG) systems. Furthermore, to evaluate video analytics in ultra-long and open-world video scenarios, we introduce a new benchmark, Ava-100. This benchmark comprises 8 videos, each exceeding 10 hours in duration, along with 120 manually annotated, diverse, and complex question-answer pairs. On Ava-100, Ava achieves top-tier performance with an accuracy of 75.8%.
The source code of Ava is available at https://github.com/I-ESC/Project-Ava. Ava-100 benchmark could be accessed at https://huggingface.co/datasets/iesc/Ava-100.
$\dagger$ Corresponding authors.
1 Introduction
Figure 1: Intelligence levels of video analytics systems.
@@ 1 Introduction
Video analytics [20, 7, 21, 30] has emerged as a transformative technology across a wide array of domains, such as environment monitoring, intelligent transportation systems, industrial automation, and retail monitoring. By leveraging the capabilities of deep learning (DL) models, video analytics systems are able to extract patterns, derive meaningful insights, and generate actionable information from video data, thereby facilitating more efficient and precise monitoring, detection, and response to events. The desired features of video analytics systems necessitate a variety of capabilities. As shown in Fig. 1, We categorize the intelligence levels of both current and prospective video analytics systems into five tiers, designated as L1 through L5:
@@
L1 systems for specific classification, segmentation, and detection using models e.g., ResNet [17] and EfficientDet [38] to extract spatial information from video data, including object classes and bounding boxes, etc [20, 7, 21].
@@
L2 systems go beyond spatial information extraction by enabling causal event detection and analytics, i.e., identifying short-term events. They use models like C3D [41] and ActionFormer [49] to detect and localize events (e.g., actions, activities, anomalies) through spatiotemporal modeling.
L3 systems advance beyond L2’s spatiotemporal detection by integrating neural language processing (NLP) capabilities. Leveraging models such as CLIPBERT [23], these systems are able to interpret and respond to natural language queries e.g., "What animals appeared in the videos?" rather than simply detecting (manually) predefined events. Although query handling remains confined to specific domains, L3 systems significantly enhance user interaction and accessibility in video analytics.
@@
Despite significant advancements, current video analytics systems [21, 7, 52, 31, 27, 37, 51, 20] primarily focus on L1 to L3 levels. These systems are designed for closed-end analytics, relying on specialized models for specific tasks, which limits their flexibility and adaptability. Consequently, we envision L4 and L5 systems to enable open-end analytics.
@@
L4 systems represent a significant leap forward by enabling open-ended video comprehension, reasoning, and analytics. These systems support general-purpose video analytics, effectively processing natural language queries and generating contextually appropriate responses, while also facilitating complex, long-term spatiotemporal reasoning. For instance, L4 systems can address inquiries such as: "What abnormal events occurred in the past ten hours?", "What caused the person to fall?", or "How did the animals behave after appearing on camera?"
@@
L5 systems advance beyond L4 by autonomously integrating external public and domain-specific knowledge sources to uncover both explicit and implicit relationships between video content and broader world knowledge. This capability fulfills the ultimate objective of video analytics: deriving profound insights and delivering actionable, automated solutions.
@@
In this paper, we delve into the development of L4 video analytics systems, leveraging the transformative potential of vision language models (VLMs). By combining vision and language understanding, VLMs enable generalized visual detection and advanced video comprehension, including causal reasoning, key information retrieval, and human-interpretable explanations. The integration of perception, reasoning, and interaction makes VLMs highly adaptable, positioning them as a key foundation for L4 systems.
@@
However, integrating VLMs into video analytics poses significant challenges, primarily due to the limited context window of current VLMs compared to the extensive duration of video sources in typical video analytics scenarios. While L1 to L3 systems handle spatial and short-term causal event detection by processing frames independently or within small sliding windows (e.g., a few seconds), L4 systems require collective analysis of related frames for long-term causal detection, summarization, and reasoning. Current VLMs, like QwenVL [42], can process up to 768 frames, covering minutes or hours of video. However, video analytics often involves much longer sources, spanning hundreds of hours or continuous streams, far beyond the capabilities of existing VLMs.
@@
Recent studies [35, 36, 44] that attempt to extend the context window of VLMs remain inadequate for processing video sources spanning hundreds of hours. Retrieval-augmented generation (RAG) frameworks [12, 16, 13, 34] aim to address similar limitations by first retrieving relevant frames from massive contents and then generating final answers. However, these approaches still face significant challenges in handling the video modality (as detailed in §7), leading to both reduced analytics accuracy and substantial computational overhead.
@@
In this paper, we present Ava, a system that integrates VLMs into video analytics to enable L4 capabilities. The core innovation of Ava is its efficient indexing mechanism, designed to handle extremely long video sources or unlimited video streams, and by leveraging the index, Ava effectively retrieves relevant information for a given query and generates accurate, robust responses. Specifically, Ava introduces the following key features that distinguish it from existing systems: 1) analyzing extremely long videos, spanning hundreds of hours or even unlimited video streams; 2) supporting near-real-time (e.g., at more than 1 FPS) index construction and analytics; and 3) handling diverse and complex queries including temporal grounding and reasoning, summarization, event and entity understanding, and key information retrieval.
@@
Particularly, Ava introduces two key components: near-real-time index construction and agentic retrieval and generation. During the index construction phase (§4), we propose event knowledge graphs (EKGs) as an indexing mechanism for video analytics. Unlike traditional knowledge graphs (KGs) used in text-based RAG systems [12, 16, 34], EKGs represent a flow of insightful events, effectively capturing video dynamics and temporal consistency. Using a small VLM, such as Qwen2.5-VL-7B, Ava extracts information to construct the EKG. To optimize this process, we introduce techniques that enable near-real-time index construction, achieving more than 5 FPS on typical edge servers equipped with 2 $\times$ 4090 GPUs.
@@
In the retrieval and generation phase (§5), Ava employs an agentic search mechanism instead of directly retrieving information from the constructed index. This approach allows Ava to proactively retrieve more relevant information by utilizing contextual hints captured within the EKG, enables Ava to handle complex queries, including summarization and multi-hop reasoning etc. Furthermore, we introduce techniques to enhance Ava to robustly generate answers based on the retrieved information.
@@
We evaluate Ava on multiple public video understanding benchmarks, including LVBench [43] and VideoMME-Long [14]. These benchmarks collectively comprise approximately 400 videos and 2,500 questions. We compare Ava against a diverse range of baselines, including state-of-the-art (SOTA) VLMs such as GPT-4o [5], Gemini-1.5-pro [39], Phi-4-Multimodal [2], Qwen2.5-VL-7B [42], InternVL2.5-8B [9], and LLaVA-Video-7B [53], as well as typical video-RAG systems such as VideoTree [47], VideoAgent [45], DrVideo [29], and VCA [48]. On both benchmarks, Ava establishes new SOTA performance, achieving 62.3% on LVBench and 64.1% on VideoMME-Long, respectively, significantly surpassing baselines by up to 16.9% and 5.2%.
@@
In addition to the public video understanding benchmarks, we introduce a new benchmark, Ava-100, specifically designed to evaluate L4 video analytics tasks. Ava-100 comprises 8 ultra-long videos, each exceeding 10 hours in duration, and includes a total of 120 manually annotated questions and answers. The videos are carefully curated from typical video analytics scenarios, and the questions cover multiple key categories relevant to video analytics systems. Evaluation results show that Ava achieves significantly better performance on Ava-100 compared to various baselines, with improvements of approximately 20.8%. In summary, we make the following contributions in this paper:
@@
  • We propose Ava, the first L4 video analytics system powered by VLMs, to the best of our knowledge.
@@
  • Ava introduces near-real-time index construction and agentic retrieval and generation, along with innovative techniques that enable key features for L4 video analytics, including open-ended analytics on extremely long video sources in near-real-time.
@@
  • We evaluate Ava on two widely-used video understanding benchmarks, achieving SOTA performance with 62.3% on LVBench and 64.1% on VideoMME-Long.
@@
  • Furthermore, we present Ava-100, a benchmark specifically designed for L4 video analytics systems, where Ava demonstrates significant improvements, outperforming baselines by approximately 20.8%.
2 Related Work and Motivation
2.1 Video Analytics System and VLMs
@@
The field of video analytics has seen significant advancements in recent years [21, 7, 52, 31, 27, 37, 51, 19]. Leveraging emerging deep learning models, modern video analytics systems can extract insightful information, such as object locations or counts, from video streams processed on edge or cloud servers.
@@
Existing video analytics systems predominantly support closed-end analytics (L1 to L3 systems as mentioned in §1), often relying on shallow neural networks to extract predefined, task-specific, and constrained information. For instance, Remix [20] leverages fine-tuned EfficientDet [38] to generate bounding boxes for pedestrians. Consequently, the flexibility and adaptability of video analytics systems are fundamentally limited by the inherent constraints of the specific models they employ.
@@
Recently, VLMs like GPT-4o [5], Gemini [39, 40], QwenVL [42, 6] and Phi [2, 3, 4], have demonstrated their transformative potential in video analytics tasks. By leveraging the extensive world knowledge embedded in large language models (LLMs), VLMs not only achieve generalized visual grounding but also exhibit advanced video comprehension capabilities, such as zero-shot temporal and spatial reasoning, contextual retrieval, and semantic understanding. More importantly, VLMs enable natural language interaction, allowing users to dynamically query video content. This makes them particularly well-suited for addressing diverse and unstructured open-ended analytics in real-world scenarios, towards L4 video analytics systems.
@@
However, adopting VLMs in video analytics systems is far from straightforward. Existing L1 to L3 video analytics systems, which rely on traditional DNNs, typically process each video frame independently. In contrast, VLMs require related frames to be processed collectively to infer causal relationships and temporal dependencies across frames. This shift introduces significant complexity, as existing VLMs are generally capable of handling only minute-level or sub-hour-level videos due to the limited context window inherent in language models.
@@
In real-world video analytics scenarios, the scale of videos to be analyzed is often vastly larger—spanning hundreds of hours or more (e.g., monitoring wildlife behavior over an entire month, as illustrated in Fig. 3). This creates a fundamental gap between the capabilities of VLMs and the demands of video analytics systems, as the limited context window length of VLMs directly restricts their ability to process videos of such extensive durations effectively.
Short (1.4 minutes) Medium (9.7 minutes) Long (39.7 minutes)
@@
Table 1: Only a small portion of the frames are necessary to answer each particular question across the short, medium, and long video subsets of VideoMME [14] benchmark using Qwen2-VL.
2.2 Long Video Understandings
Recent research also increasingly focuses on enabling long video understandings [35, 44]. Current autoregressive language models inherently have the constrained context window length, therefore efforts have been directed toward reducing the redundancy in video inputs to facilitate the processing of extended video durations. For instance, LongVU [35] and AdaRETAKE [44] introduce dynamic compression mechanisms that prioritize video content based on its relevance, selectively retaining frames or regions most pertinent to downstream language tasks. Similarly, NVILA [25] addresses the efficiency-accuracy trade-off by optimizing sampling strategies and resolution to fit within limited token budgets.
While these approaches have succeeded in increasing the number of frames that models can process and mitigating the constraints of context windows to some extent, they fall short of achieving a fundamental breakthrough. Given that SOTA LLMs typically support context lengths ranging from 128K to 1M tokens, most existing approaches are thereby restricted to processing video segments of at most one hour in duration [32], which is inadequate for the requirements of video analytics. Furthermore, as video length increases, the associated inference cost rises proportionally, thereby compounding the scalability challenges inherent to these systems.
2.3 Retrieval Augmented Generation
Figure 2: The system overview of Ava.